A projectile shot at an angle of 60° above the horizontal strikes a building 25m away at a point 16m above the point of projection. Find the magnitude and direction of the velocity of the projectile as it strikes the building
1
Expert's answer
2020-06-21T20:15:23-0400

Let the x axis be directed from the point of the start of the motion towards the base of the building and the y axis be perpendicular to the x axis and directed vertically.


The projection of the initial velocity onto x axis is "v_x = v_0\cos\alpha" and onto y axis is "v_y=v_0\sin\alpha - gt." Therefore, x coordinate will be "x=v_0\cos\alpha t, \; y = v_0\sin\alpha t - \dfrac{gt^2}{2}."

We may write the system

"\begin{cases}
x=v_0\cos\alpha t, \\
y = v_0\sin\alpha t - \dfrac{gt^2}{2}.
\end{cases}
\;\;\; 

\begin{cases}
25=v_0\cos60^\circ t, \\
16 = v_0\sin60^\circ t - \dfrac{9.8t^2}{2}.
\end{cases}
\;\;\;
\begin{cases}
t=\dfrac{25}{v_0\cos60^\circ}, \\
16 = 25\tan60^\circ - \dfrac{9.8\cdot\dfrac{25^2}{v_0^2\cos^260^\circ}}{2}.
\end{cases}"

Therefore, "v_0 = 21.2\,\mathrm{m/s}, \; t= 2.36\,\mathrm{s}." The velocity components will be

"v_x = v_0\cos\alpha = 10.6\,\mathrm{m/s}, \; v_y = v_0\sin\alpha - gt = -4.77\,\mathrm{m/s}." So the total velocity will be "\sqrt{v_x^2+v_y^2} = 11.6\,\mathrm{m/s}" . The angle between the direction of x axis and the velocity is "\tan\beta = \dfrac{v_y}{v_x} = -0.45, \beta = -24^\circ."